OpenStudy (anonymous):

10x=-80, x=?, solve using multiplication principle

OpenStudy (anonymous):

The idea behind this is that every element of the field has a multiplicative inverse.

OpenStudy (anonymous):

\[10x\cdot10^{-1} = 10x\cdot\frac{1}{10} = -80 \cdot \frac{1}{10}\]

OpenStudy (anonymous):

Now applying commutativity of multiplication we have \[\frac{10}{10} \cdot x = \frac{-80}{10}\]\[x = -8\]

OpenStudy (anonymous):

Is that the so called "multiplication principle" math teachers talk about?

OpenStudy (anonymous):

Or you could note that 10 *x = 10*-8...